OpenStudy (anonymous):

The number of mole of a molecule is defined as its mass per molar mass. Using # mol = n, mass = m and molar mass = M and given that n = M m , rearrange this expression for (a) M = ? (b) m = ? (c) If n = 0.0021 mol and M = 44 g/mol; what is the value of m?

OpenStudy (chmvijay):

n=m/M M=m/n m=M*n